### v1

# On two uniform exponents of approximation related to Wirsing’s problem

Abstract.
We aim to fill a gap in the proof of an inequality relating two exponents of uniform Diophantine approximation stated in a paper by Bugeaud. We succeed to verify the inequality in several instances, in particular for small dimension. Moreover, we provide counterexamples to generalizations, which contrasts the case of ordinary approximation where such phenomena do not occur. Our results contribute to the understanding of the discrepancy between small absolute values of a polynomial at a given real number and approximation to the number by algebraic numbers of absolutely bounded degree, a fundamental issue in the famous problem of Wirsing and variants.
Middle East Technical University, Northern Cyprus Campus, Kalkanli, Güzelyurt
johannes@metu.edu.tr ; jschleischitz@outlook.com
Keywords: Wirsing’s problem, exponents of Diophantine approximation
Math Subject Classification 2010: 11J13, 11J82
1. Introduction
1.1. Some exponents of approximation and basic properties
Let $n\geq 1$ be an integer and $\xi$ be a transcendental real number. For a polynomial $P$, by its height $H_{P}$ we mean the maximum modulus of its coefficients. For a parameter $X>1$, we call the non-zero integer polynomial $P_{X}$ of degree at most $n$ and height at most $X$ the best approximation polynomial for $n,\xi$ up to parameter $X$, if $P=P_{X}$ minimizes $|P(\xi)|$ among all such integer polynomials $P$. This is defined uniquely up to sign since $\xi$ is transcendental. Similarly, write $\alpha_{X}$ for the algebraic number of degree at most $n$ and height at most $X$ that minimizes $|\xi-\alpha_{X}|$. Here the height of $\alpha$, let us denote it $H(\alpha_{X})$, is simply the height of its minimal polynomial over the integers with coprime coefficients. Define the local exponents of approximation at parameter $X>1$ by
$w(n,\xi,X)=-\frac{\log|P_{X}(\xi)|}{\log X},\qquad w^{\ast}(n,\xi,X)=-\frac{\log(H(\alpha_{X})\cdot|\alpha_{X}-\xi|)}{\log X}$
and let us derive the classical exponents of approximation
(1) $w_{n}(\xi)=\limsup_{X\to\infty}w(n,\xi,X),\qquad w_{n}^{\ast}(\xi)=\limsup_{X\to\infty}w^{\ast}(n,\xi,X)$
and
(2) $\widehat{w}_{n}(\xi)=\liminf_{X\to\infty}w(n,\xi,X),\qquad\widehat{w}_{n}^{\ast}(\xi)=\liminf_{X\to\infty}w^{\ast}(n,\xi,X).$
The ordinary exponents (without “hat”) may take the formal value +$\infty$+ ∞, on the other hand the uniform exponents (with “hat”) are bounded independently of $\xi$ in terms of $n$, see Davenport and Schmidt [10] and also [8]. The exponents $w_{n}(\xi)$ date back to Mahler, the exponents $w_{n}^{\ast}(\xi)$ were introduced by Wirsing [22] in his famous paper on approximation to real numbers by algebraic numbers of bounded degree. The uniform exponents $\widehat{w}_{n}(\xi),\widehat{w}_{n}^{\ast}(\xi)$ were introduced by Bugeaud and Laurent [7]. It is easily seen that all four types of exponents from (1), (2) form non-decreasing sequences in $n$, in particular we have
(3) $w_{1}(\xi)\leq w_{2}(\xi)\leq\cdots,\qquad\widehat{w}_{1}^{\ast}(\xi)\leq\widehat{w}_{2}^{\ast}(\xi)\leq\cdots.$
Several other relations are known, most notably Dirichlet’s Theorem implies
(4) $w_{n}(\xi)\geq\widehat{w}_{n}(\xi)\geq n,\qquad n\geq 1.$
A famous open problem of Wirsing [22] asks if a strengthening of (4) of the form $w_{n}^{\ast}(\xi)\geq n$ holds as well. As a rather short argument (see Proposition 4.7 below) yields for any $\xi$ the estimates
(5) $\widehat{w}_{n}^{\ast}(\xi)\leq\widehat{w}_{n}(\xi),\qquad w_{n}^{\ast}(\xi)\leq w_{n}(\xi),$
the conjectural inequality is indeed stronger, and verified for $n\in\{1,2\}$ only so far [9]. It was already noticed by Wirsing that for any $n$ and almost all real numbers in the sense of Lebesgue measure $w_{n}^{\ast}(\xi)=n$, in fact $\widehat{w}_{n}^{\ast}(\xi)=n$ as well as a consequence of (28) below from [7], and Sprindžuk’s result [21]. However, for general $\xi$ the best known lower bounds for $w_{n}^{\ast}(\xi)$ and large $n$ are of order $n/\sqrt{3}$ from [1]. Since $\widehat{w}_{1}(\xi)=\widehat{w}_{1}^{\ast}(\xi)$ for any transcendental real $\xi$, combining (3), (4) we get
(6) $w_{n}^{\ast}(\xi)\geq\widehat{w}_{n}^{\ast}(\xi)\geq 1,\qquad n\geq 1.$
Let us further define
$\kappa(n,\xi,X):=w(n,\xi,X)-w^{\ast}(n,\xi,X),$
and accordingly
$\underline{\kappa}_{n}(\xi):=\liminf_{X\to\infty}\kappa(n,\xi,X),\quad\overline{\kappa}_{n}(\xi):=\limsup_{X\to\infty}\kappa(n,\xi,X).$
These exponents have not been explicitly defined or studied before. Note that
(7) $\overline{\kappa}_{n}(\xi)\geq\max\{w_{n}(\xi)-w_{n}^{\ast}(\xi),\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\}$
and conversely
(8) $\underline{\kappa}_{n}(\xi)\leq\min\{w_{n}(\xi)-w_{n}^{\ast}(\xi),\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\}$
as the according estimates are true for general pairs of real functions $f(X),g(X)$ in place of $w(n,\xi,X)$ and $w^{\ast}(n,\xi,X)$. Moreover a generalization of both claims of (5) of the form
(9) $\underline{\kappa}_{n}(\xi)\geq 0,$
holds, by the same Proposition 4.7. Note however that there is no reason for the quantities $\underline{\kappa}_{n}(\xi),\overline{\kappa}_{n}(\xi)$ to be monotonic in $n$, in particular an analogue of (3) does not hold. We believe that these exponents carry important information on the discrepancy between small values of polynomials and approximation to real numbers by algebraic numbers, closely related to Wirsing’s problem and variants, and deserve to be studied in detail.
1.2. An open problem and related questions
Let us return to the classical exponents. As stated in [5], conversely to (5) we have
(10) $w_{n}(\xi)-w_{n}^{\ast}(\xi)\leq n-1.$
This follows from combination of Lemma 4.1 and Lemma 4.2 below. The sharpness of (10) is open in general, for partial results see [4]. It was further stated in [6] that we similarly have
(11) $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq n-1.$
The latter estimate is sharp, as it was noticed in [6] that any Liouville number $\xi$ (i.e. $\xi$ is transcendental and $w_{1}(\xi)=\infty)$ satisfies for any $n\geq 1$ the identities
(12) $\widehat{w}_{n}(\xi)=n,\qquad\widehat{w}_{n}^{\ast}(\xi)=1.$
The left identity stems from [16], see Theorem 4.6 below. In fact, it was shown in [6] that the interval $[n-2+\frac{1}{n},n-1]$ is contained in the spectrum of the difference $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)$, and upon assuming some natural conjecture to be true, the spectrum contains $[0,n-1]$.
However, there seems to be a considerable gap in the sketched proof of (11). Indeed, the proof is based on Lemma 4.2 below, which however only holds for separable $P$. In contrast to (10) for the ordinary exponents, where this problem can be overcome rather easily by Lemma 4.1 due to Wirsing below, for uniform exponents it is a priori unclear why one can restrict to this case. While at first sight this may appear to be a rather minor technicality, it turns out to cause great problems, as evidenced by the results in § 3.2 below.
The purpose of this note is twofold: Firstly, we want to address the very problem of rigorously settling (11). This will be the content of § 2. Secondly, in § 3, we study the related parametric functions $\kappa(n,\xi,X)$ and their extremal values $\underline{\kappa}_{n}(\xi),\overline{\kappa}_{n}(\xi)$. We once again stress that understanding these quantities is at the core of Wirsing’s famous problem discussed in § 1.1.
To this end, we again want to emphasize that the delicacy of proving (11) is reflected in our results in § 3.2 below, where we will provide counterexamples to stronger claims involving the exponents $\kappa$ and their extremal values. In particular, Theorem 3.3 below provides a counterexample of a parametric inequality in the spirit of (9) that if true would generalize both (10), (11). This illustrates that indeed one has to be very careful with non separable polynomials.
2. Results towards a proof of (11)
Despite the examples from § 3.2 mentioned above that suggest otherwise, we still believe that (11) is true for all $n$. While we were unable to prove (11) in full generality, we provide several weaker claims. First we verify it for small $n$.
Theorem 2.1.
Claim (11) holds for $n\leq 5$ and any transcendental real number $\xi$.
While the conclusion for $n\leq 3$ is short by combining results from [5, 16], the cases $n=4,n=5$ require a more sophisticated approach. With more effort it is likely that the range for $n$ can be extended by the underlying arguments of the latter cases. However we did not intensify our investigations in this direction, and are doubtful that the strategy is sufficient for a proof for general $n$ without further new ideas.
Next we want to state criteria implying (11) for arbitrary $n$. In this context we want to define another classical exponent of approximation. Let the ordinary exponent of simultaneous approximation denoted by $\lambda_{n}(\xi)$ be the supremum of $\lambda$ so that
$\max_{1\leq j\leq n}|q-\xi^{j}p_{j}|\leq q^{-\lambda}$
has infinitely many solutions in integer vectors $q,p_{1},\ldots,p_{n}$.
Theorem 2.2.
Let an integer $n\geq 1$ and a transcendental real number $\xi$ satisfy any of the following conditions
  • (i)
    $\widehat{w}_{n}(\xi)=n$
  • (ii)
    $\widehat{w}_{n}(\xi)>2n-3$
  • (iii)
    $\widehat{w}_{n}^{\ast}(\xi)\geq n-2$
  • (iv)
    $\lambda_{n}(\xi)\in[\frac{1}{n},\frac{1}{n-2}]\cup(1,\infty]$
  • (v)
    $w_{1}(\xi)\geq n$
  • (vi)
    $w_{n}(\xi)\leq\frac{2n-1+\sqrt{4n-3}}{2}$
Then (11) holds.
Presumably (ii) can only happen for small $n$. It is known that $\widehat{w}_{n}(\xi)\leq 2n-2$ for $n\geq 10$, and slightly weaker bounds are settled for $n\leq 9$, see [17]. Notably, for $n\geq 3$ it is in fact not known if counterexamples to (i) exist. By (4), one of the conditions (i), (ii) in particular applies to any $\xi$ if $n\leq 3$ (and (iii) does as well), so part of Theorem 2.1 is directly implied. However, the claim for $n=2$ from Theorem 2.1 is in fact used in the proof of this case.
A corollary of part (ii) is the following unconditional but weaker bound that may be of some interest for small $n$ especially.
Theorem 2.3.
For any integer $n\geq 3$ and any transcendental real number $\xi$ we have
$\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq 2n-4.$
Proof.
If $\widehat{w}_{n}(\xi)>2n-3$ then the stronger bound $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq n-1\leq 2n-4$ for $n\geq 3$ applies by Theorem 2.2, (ii). If otherwise $\widehat{w}_{n}(\xi)\leq 2n-3$ then (6) implies $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq\widehat{w}_{n}(\xi)-1\leq 2n-4$ as well. ∎
This slightly improves on the bound $22n-$ 2 that is immediate from Davenport and Schmidt [10], in fact $32n-$ 3 for $n\geq 10$ and slightly weaker claims for smaller $n$ follow from [8]. It would be desirable to settle a bound stronger than $2n-O(1)$.
3. On the exponents $\kappa(n,\xi,X)$
3.1. On $\overline{\kappa}_{n}$
Our next theorem gives another upper estimate for the difference $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)$, in fact for the quantity $\overline{\kappa}_{n}(\xi)$. However another exponent of approximation occurs in the formula. Write $\lfloor x\rfloor$ for the largest integer less or equal than $x$.
Theorem 3.1.
Let $n\geq 2$ be an integer and derive
$k=\left\lfloor\frac{n}{2}\right\rfloor.$
Then for any real number $\xi$ we have
(13) $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq\max\{n-1,w_{k}(\xi)\}.$
In fact the stronger inequality
(14) $\overline{\kappa}_{n}(\xi)\leq\max\{n-1,w_{k}(\xi)\}$
holds.
By (7), estimate (14) indeed implies (13). The claim shows that the desired estimate (11) holds if $w_{k}(\xi)\leq n-1$. The classical formula of Bernik [3]
(15) $\dim_{H}(\{\xi:w_{m}(\xi)\geq\lambda\})=\frac{m+1}{\lambda+1},\qquad m\geq 1,\;\lambda\in[m,\infty],$
where $\dim_{H}$ denotes Hausdorff dimension, applied for $m=k$ in particular implies
Corollary 3.2.
Given an integer $n\geq 2$, define the set of counterexamples to (11) as
$\Omega_{n}:=\{\xi\in\mathbb{R}:\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)>n-1\}$
and let similarly
$\Gamma_{n}:=\{\xi\in\mathbb{R}:\overline{\kappa}_{n}(\xi)>n-1\}\supseteq\Omega_{n}.$
Then
$\dim_{H}(\Omega_{n})\leq\dim_{H}(\Gamma_{n})\leq\frac{1}{2}+o(1),\qquad n\to\infty.$
By Theorem 2.1, the estimates for $\Omega_{n}$ are only relevant for $n\geq 6$. While the bound in Corollary 3.2 seems not particularly strong, it is not clear if there is any other easy argument available that improves on it, even for $\Omega_{n}$. Indeed, it seems no improvement via Theorem 2.2 can be obtained when combining it with known metrical theory for classical exponents, and only (i), (iii) have potential to lead to such an improvement if the metrical theory is developed well enough (for (iv) this is excluded for $n\geq 6$ by a metrical result from [2]). In this context note that the Hausdorff dimension of counterexamples to (iii) is at least $1/4-o(1)$ as $n\to\infty$, as a consequence of [19, Theorem 2.5] (in fact of its proof) and (15).
For a reverse positive lower bound for the Hausdorff dimension of $\Gamma_{n}$, see Corollary 3.4 below. For $n\in\{2,3\}$ we will improve the bounds resulting from Corollary 3.2 in Corollary 6.2 below.
3.2. Counterexamples to stronger versions of (11)
The following example shows that a parametric version of (11) does not hold in general, so that $w_{k}(\xi)$ cannot be dropped in (14) of Theorem 3.1. Recall that a Liouville number is a transcendental real number satisfying $w_{1}(\xi)=\infty$.
Theorem 3.3.
Let $n\geq 2$ be an integer and $\xi$ be a transcendental real number. Let $2\leq k\leq n$ be another integer. If $w_{1}(\xi)>n+k-1$ then
(16) $\overline{\kappa}_{n}(\xi)\geq\left(1-\frac{1}{k}\right)\cdot w_{1}(\xi).$
Thus, if $w_{1}(\xi)>2n-1$, then
(17) $\overline{\kappa}_{n}(\xi)\geq\left(1-\frac{1}{n}\right)\cdot w_{1}(\xi).$
In particular, if $\xi$ is a Liouville number then
(18) $\overline{\kappa}_{n}(\xi)=\infty.$
The proof is based on Liouville’s inequality Theorem 4.4 below on the minimum distance between algebraic numbers. As another application of Theorem 3.3, we get
Corollary 3.4.
We have
(19) $\dim_{H}(\Gamma_{n})\geq\dim_{H}(\Gamma_{n}\setminus\Omega_{n})\geq\frac{2-o(1)}{n+\sqrt{n}},\qquad n\to\infty,$
with $\Gamma_{n}$ defined in Corollary 3.2. Moreover we always have $\dim_{H}(\Gamma_{n}\setminus\Omega_{n})\geq 1/n$.
Proof.
Let $k$ be the smallest integer with $k^{2}-k+1>n$, of order $\sqrt{n}$. Then the right hand side of (16) exceeds $n-1$ as soon as $w_{1}(\xi)>n+k-1$, which also justifies application of (16). Then we can further use (15) for $m=1$ and any $\lambda>n+k-1$, which gives the lower bound of (19) for the set $\Gamma_{n}$. Finally the hypothesis implies (11) for all $\xi$ within our set from (15) by Theorem 2.2, (v), hence the estimate is preserved when removing $\Omega_{n}$. The bound $2/(2n)=1/n$ follows similarly from (17) with $\lambda=2n-1+\varepsilon$ via (15). ∎
Similarly parametric variants of (19) for estimating from below
(20) $\dim_{H}(\{\xi\in\mathbb{R}:\overline{\kappa}_{n}(\xi)>\lambda\}),\qquad\lambda\geq\lambda_{0}(n)$
with large enough $\lambda_{0}(n)$ can be obtained. Moreover, variants of Theorem 3.3 involving $w_{\ell}(\xi)$ for any $\ell\leq n/2$ can be derived, in particular we have a generalization of (18) of the form $\overline{\kappa}_{n}(\xi)=\infty$ as soon as $\xi$ is a $U_{\ell}$-number of index $\ell\leq n/2$ in Mahler’s classification of real numbers.
As indicated in § 1.1, inseparable polynomials cause problems when trying to prove (11), the deeper reason being that we cannot apply Lemma 4.2 and Lemma 6.1 below. This motivates to define the separable local exponent $w_{sep}(n,\xi,X)$ like $w(n,\xi,X)$ but restricting to separable polynomials $P$. Let us derive the uniform exponent
$\widehat{w}_{n,sep}(\xi):=\liminf_{X\to\infty}w_{sep}(n,\xi,X).$
Hence, in order to prove (11), one may try to settle for any real $\xi$ the identity
(21) $\widehat{w}_{n,sep}(\xi)=\widehat{w}_{n}(\xi).$
Indeed, if (21) holds for some $n,\xi$, then (11) holds for this pair via
$\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)$ $=\widehat{w}_{n,sep}(\xi)-\widehat{w}_{n}^{\ast}(\xi)=\liminf_{X\to\infty}w_{sep}(n,\xi,X)-\liminf_{X\to\infty}w^{\ast}(n,\xi,X)$
$\leq\limsup_{X\to\infty}(w_{sep}(n,\xi,X)-w^{\ast}(n,\xi,X))\leq n-1,$
where for the last inequality we used Lemma 4.2 below. For completeness, define similarly $w_{irr}(n,\xi,X)$ and $\widehat{w}_{n,irr}(\xi)$ when restricting to irreducible polynomials. It is clear that
$w_{irr}(n,\xi,X)\leq w_{sep}(n,\xi,X)\leq w(n,\xi,X),$
hence for any $n,\xi$ we have
$\widehat{w}_{n,irr}(\xi)\leq\widehat{w}_{n,sep}(\xi)\leq\widehat{w}_{n}(\xi).$
We remark that one may accordingly define ordinary exponents $w_{n,irr}(\xi),w_{n,sep}(\xi)$ as well via upper limits, however they turn out to be equal to $w_{n}(\xi)$ by Lemma 4.1 below.
However, at least for $n=2$ we can prove that (21) is in general false as well, so this avenue is not possible.
Theorem 3.5.
Assume $\xi$ satisfies $w_{1}(\xi)>3$. Then
(22) $\widehat{w}_{2,sep}(\xi)\leq 1+\frac{2}{w_{1}(\xi)-1}<2\leq\widehat{w}_{2}(\xi).$
In particular, for any Liouville number $\xi$ we have
$\widehat{w}_{2,irr}(\xi)=\widehat{w}_{2,sep}(\xi)=1.$
Remark 1.
A slightly stronger estimate for $\widehat{w}_{2,irr}(\xi)$ upon $w_{1}(\xi)\in(3,\infty)$ can be deduced from the method. We further remark that as shown in [18], the exponent $\widehat{w}_{=n,irr}(\xi)$ (denoted just $\widehat{w}_{=n}(\xi)$ in [18]) with polynomials irreducible of degree precisely $n$ equals 00 for any Liouville number and any $n\geq 2$. The proof is similar to our proof of Theorem 3.5 below, based on Minkowski’s Second Convex Body Theorem. In exact degree the associated ordinary exponents $w_{=n,irr}(\xi)$ become more interesting as well. In particular, it is open for $n>7$ if it is always at least $n$ (for $n\leq 7$ this was settled in [20, Theorem 1.3]).
We believe there is equality in the most left inequality of (22). Denoting $spec(.)\subseteq\mathbb{R}$ the set of values taken by an exponent when inserting all real numbers, this would imply
(23) $spec(\widehat{w}_{2,sep})=[1,2]\cup spec(\widehat{w}_{2}),\qquad spec(\widehat{w}_{2}-\widehat{w}_{2,sep})=[0,1].$
The spectrum of $\widehat{w}_{2}$ is known to be contained and dense in $[2,(3+\sqrt{5})/2]$. The inclusion of $[1,2]$ in the left identity of (23) follows directly from equality in (22). On the other hand, if $\widehat{w}_{2}(\xi)>2$ for some $\xi$ then it follows as in [18] that all best approximations are irreducible of degree exactly two, which implies $\widehat{w}_{2,sep}(\xi)=\widehat{w}_{2}(\xi)$. This argument shows that the spectrum of $\widehat{w}_{2,spec}(\xi)$ contains the spectrum of $\widehat{w}_{2}(\xi)$ as well, more precisely unconditionally we have
$spec(\widehat{w}_{2})\cap(2,\infty)=spec(\widehat{w}_{2,sep})\cap(2,\infty).$
The point $\{2\}$ can be added by considering generic numbers with $w_{2}(\xi)=2$, see Theorem 3.7 below. Hence the left identity of (23) holds upon our assumption. The right identity of (23) uses additionally that $w_{1}(\xi)\geq 3>2$ implies $\widehat{w}_{2}(\xi)=2$ by Theorem 4.6 below.
We further believe that for larger $n$ again Liouville numbers will provide counterexamples to (21), however the proof will be more involved so we do not attempt to make it rigorous.
3.3. Some remarks on $\underline{\kappa}_{n}$ and the limit of $\kappa$
Let us study $\underline{\kappa}_{n}(\xi)$. By (9), (10), (8) we get
(24) $10\leq\underline{\kappa}_{n}(\xi)\leq w_{n}(\xi)-w_{n}^{\ast}(\xi)\leq n-$ 1 .
This motivates the following
Problem 1.
Can we improve the bound $n-1$ for $\underline{\kappa}_{n}$ from (24) for $n\geq 2$ and all $\xi$? Is it in fact always 00?
Let us first consider $n=2$. Then, complementary to the bound $n-1=1$ in (24), we get the following estimates in terms of the uniform exponent.
Theorem 3.6.
We have
(25) $\underline{\kappa}_{2}(\xi)\leq\frac{w_{2}(\xi)}{\widehat{w}_{2}(\xi)-1}-\widehat{w}_{2}(\xi)$
and
(26) $\underline{\kappa}_{2}(\xi)\leq\frac{1}{\widehat{w}_{2}(\xi)-1}.$
The latter (26) is a refinement of (24) by (4). Note that the bound in (25) is always non-negative [11]. We enclose a few observations on Problem 1 for general $n$:
  • If $\underline{\kappa}_{n}(\xi)=0$ for all $\xi$, this would imply Wirsing’s problem has an affirmative answer for given $n$. In fact $\underline{\kappa}_{n}(\xi)=0$ implies $w_{n}^{\ast}(\xi)\geq\widehat{w}_{n}(\xi)\geq n$. While this is audacious to conjecture in general, for $n=2$, Theorem 3.6 and the results from [9, 14] it is based on indicate that it may be true.
  • Constructing $\xi$ with $\underline{\kappa}_{n}(\xi)>0$ does not seem easy either, as all best approximations of large norm must have untypically small derivative. Moreover it probably forces $\widehat{w}_{n}(\xi)>n$, see Conjecture 1 below.
  • An analogue of Jarník’s estimate on the logarithmic ratio of best approximation norms, as in the proof of Theorem 3.6 below, for higher dimension would generalize Theorem 3.6. However, this seems to this date unknown. While the results from [13] suggest such an effective estimate, it seems not settled. Nevertheless, in view of Conjecture 1 below, this suggests that $n-1$ can be improved. Moreover, it seems that Wirsing’s method and results [22] do not immediately imply anything better than $n-1$ either. For example if $w_{n}(\xi)=2n-1$ (or larger).
The next metrical result on the limit of $\kappa$ is an easy consequence of a claim from [7].
Theorem 3.7.
Let $n\geq 2$ be an integer and $\xi$ be a transcendental real number. If
$w_{n}(\xi)=n$
then
(27) $\lim_{X\to\infty}\kappa(n,\xi,X)=0.$
In particular we have (27) for almost all $\xi$ with respect to Lebesgue measure.
Proof.
We use the estimate of Bugeaud and Laurent [7, Theorem 2.1]
(28) $\widehat{w}_{n}^{\ast}(\xi)\geq\frac{w_{n}(\xi)}{w_{n}(\xi)-n+1}.$
It follows that
$\overline{\kappa}_{n}(\xi)$ $=\limsup_{X\to\infty}(w(n,\xi,X)-w^{\ast}(n,\xi,X))$
$\leq\limsup_{X\to\infty}w(n,\xi,X)-\liminf_{X\to\infty}w^{\ast}(n,\xi,X)$
(29) $=w_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq w_{n}(\xi)-\frac{w_{n}(\xi)}{w_{n}(\xi)-n+1}.$
For $w_{n}(\xi)=n$ the right hand side vanishes, thus together with (9) we conclude
$0\leq\underline{\kappa}_{n}(\xi)\leq\overline{\kappa}_{n}(\xi)\leq 0$
so the limit is 00. Finally for the metrical claim we conclude with Sprindžuk’s famous result [21]. ∎
More generally combining (3.3) and (15) enables us to estimate from above (20) for $\lambda\geq 0$, complementary to Corollary 3.2 where $\lambda=n-1$. For large $\lambda$, only small improvements compared to the bounds obtained via trivially estimating $\widehat{w}_{n}^{\ast}(\xi)\geq 1$ in place of (28) and using (15) are achieved. In particular for $\lambda=n-1$ the bound becomes significantly weaker than the one from Corollary 3.2.
The proof of Theorem 3.7 and the dual inequality $w_{n}^{\ast}(\xi)\geq\widehat{w}_{n}(\xi)/(\widehat{w}_{n}(\xi)-n+1)$ to (28) established in [7] as well suggest the following may be true.
Conjecture 1.
We have
(30) $\underline{\kappa}_{n}(\xi)\leq\widehat{w}_{n}(\xi)-\frac{\widehat{w}_{n}(\xi)}{\widehat{w}_{n}(\xi)-n+1}.$
In particular
$\widehat{w}_{n}(\xi)=n$
implies
$\underline{\kappa}_{n}(\xi)=0.$
However, the conclusion even for the special latter claim is not as straightforward as for Theorem 3.7, and we could not prove it rigorously. Note also that $\widehat{w}_{n}(\xi)=n$ does not imply the stronger claim $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)=0$, as any Liouville number $\xi$ satisfies (12). Combining (26) with our unproved estimate (30) for $n=2$ gives a conditional bound
$\underline{\kappa}_{2}(\xi)\leq\frac{1}{\sqrt{2}}.$
4. Preparatory lemmas
We need some lemmas. The first is essentially due to Wirsing [22, Hilfssatz 4].
Lemma 4.1 (Wirsing).
Let $\xi$ be a transcendental real number and $n\geq 1$. Let $\epsilon>0$. Let $P$ be an integer polynomial of degree at most $n$ and large enough height, and assume for some $w>0$ we have
$|P(\xi)|\leq H_{P}^{-w}.$
Then there exists an irreducible divisor $Q$ of $P$ such that
$|Q(\xi)|\leq H_{Q}^{-w+\epsilon}.$
We note that this easily implies (10). The next is part of Lemma A8 in Bugeaud’s book [5] attributed to Feldman.
Lemma 4.2 (Feldman).
If $P$ is a separable integer polynomial of degree at most $n$ then it has a root $\alpha$ such that
$|\xi-\alpha|\ll_{n}|P(\xi)|\cdot H_{P}^{n-2}.$
Thus without restriction to separable polynomials, the claim (11) would be obvious. The next well-known claim is the essence in the proof of Lemma 4.1 above.
Lemma 4.3 (Gelfond’s Lemma).
For any polynomials with $P=QR$ of degree at most $n$ we have
$H_{P}\asymp_{n}H_{Q}H_{R}.$
We state Liouville’s inequality, see [5, Corollary A2].
Theorem 4.4 (Liouville inequality).
If $\alpha$ is real algebraic of degree $m$ and $\beta$ is real algebraic of degree $n$ and they have different minimal polynomials then
$|\alpha-\beta|\gg_{m,n}H(\alpha)^{-n}H(\beta)^{-m}.$
The next is [8, Theorem 2.3].
Theorem 4.5 (Bugeaud, Schleischitz).
For any positive integers $m,n$ and any real transcendental number $\xi$, we have
$\min\{w_{m}(\xi),\widehat{w}_{n}(\xi)\}\leq m+n-1.$
The next is combination of [16, Theorem 1.12] and [16, Theorem 1.6].
Theorem 4.6 (Schleischitz).
Let $n\geq 1$ be an integer and $\xi$ be a transcendental real number. If $w_{1}(\xi)\geq n$ then $\widehat{w}_{n}(\xi)=n$. If $\lambda_{n}(\xi)>1$ then $w_{1}(\xi)>2n-1\geq n$, so again $\widehat{w}_{n}(\xi)=n$.
For completeness, we end this section with a well known fact although not explicitly used in proofs below, see for example [5] for a proof.
Proposition 4.7.
For any integer polynomial $P$ of degree at most $n$, any root $\alpha$ of it satisfies
$|\xi-\alpha|\gg_{n,\xi}H(\alpha)^{-1}|P(\xi)|.$
Proposition 4.7 immediately proves (5), (9).
For the proofs below recall that $P=P_{X}$ minimizes $|P(\xi)|$ among all integer polynomials of degree at most $n$ and height at most $X$. It satisfies
$|P_{X}(\xi)|\leq X^{-\widehat{w}_{n}(\xi)+o(1)},\qquad X\to\infty.$
5. Proof of Theorem 3.1
Let $\epsilon>0$. Let $X$ be any large parameter and $P=P_{X}$ of height $H_{P}\leq X$ be the according best approximation polynomial. For simplicity write $\tilde{w}=w(n,\xi,X)$. Define $\sigma=\sigma_{X}$ implicitly by
$X=H_{P}^{\sigma},\qquad\sigma=\sigma_{X}\geq 1.$
Then
$|P(\xi)|=H_{P}^{-\sigma\tilde{w}}.$
By Wirsing’s Lemma 4.1 and again assuming $X$ was chosen sufficiently large, there is an irreducible (thus separable) divisor $Q=Q_{X}$ of $P$ so that
$|Q(\xi)|=H_{Q}^{-\sigma w},\qquad w\geq\tilde{w}-\epsilon.$
Define $\tau=\tau_{X}$ implicitly via
$H_{Q}^{\tau}=X,\qquad\tau=\tau_{X}\geq 1-\epsilon.$
The lower bound is obtained from Gelfond’s Lemma 4.3 which in fact implies the stronger claim $H_{Q}\ll_{n}H_{P}\leq X$. Let $m=m_{X}=\deg Q\leq n$ and $r=n-m=\deg R$ where $P=QR$.
We may assume
(31) $m>k$
for large enough $X$. Indeed otherwise if $m\leq k$ for arbitrarily large $X$ then for some modified small $\varepsilon>0$ we get
$w_{k}(\xi)\geq w_{m}(\xi)\geq\sigma\tilde{w}-\varepsilon\geq\tilde{w}-\varepsilon.$
which further implies
$\tilde{w}\leq w_{k}(\xi)+\varepsilon\leq\max\{w_{k}(\xi),n-1\}+w^{\ast}(n,\xi,X)+\varepsilon,$
and the claim follows by subtracting $w^{\ast}(n,\xi,X)$ from both sides as $X\to\infty$ and thus $\varepsilon\to 0$.
To simplify the presentation, let us from now on consider a sequence of $P$ and assume $H_{P}\to\infty$ implying $X\to\infty$, and use $o(1)$ notation as $X\to\infty$. Similar to the proof of Theorem 3.6, Feldman’s Lemma 4.2 yields that the polynomial $Q$ has a root $\alpha$ so that
$|\alpha-\xi|\ll_{n}H(\alpha)^{-1}|Q(\xi)|H_{Q}^{m-1}=H(\alpha)^{-1}X^{-\sigma w/\tau}X^{(m-1)/\tau}=H(\alpha)^{-1}X^{-\frac{\sigma w-m+1}{\tau}}.$
Hence
(32) $w^{\ast}(n,\xi,X)\geq\frac{\sigma w-m+1}{\tau}-o(1),\qquad X\to\infty.$
On the other hand, if we assume $H_{R}\to\infty$ as $X\to\infty$, we get
$|P(\xi)|=|Q(\xi)|\cdot|R(\xi)|\geq X^{-\sigma w/\tau}\cdot H_{R}^{-w_{r}(\xi)-o(1)},\qquad X\to\infty.$
Otherwise if $H_{R}$ is bounded on a subsequence as $X\to\infty$ then by transcendence of $\xi$ we get stronger estimates from $|R(\xi)|\gg 1$. Now since by Gelfond’s Lemma 4.3
$H_{R}\asymp_{n}\frac{H_{P}}{H_{Q}}=X^{1/\sigma}\cdot X^{-1/\tau}=X^{\frac{\tau-\sigma}{\sigma\tau}}$
we infer
(33) $w(n,\xi,X)\leq\frac{\sigma w}{\tau}+w_{r}(\xi)\frac{\tau-\sigma}{\sigma\tau}+o(1),\qquad X\to\infty.$
Note that
$r=n-m\leq n-k-1\leq k$
by (31) and definition of $k$. Together with (3) this implies
$w_{r}(\xi)\leq w_{k}(\xi).$
Hence comparing the bounds (32), (33) we get
$\kappa(n,\xi,X)$ $=w(n,\xi,X)-w^{\ast}(n,\xi,X)\leq w_{r}(\xi)\frac{\tau-\sigma}{\sigma\tau}+\frac{m-1}{\tau}+o(1)$
(34) $\leq w_{k}(\xi)\frac{\tau-\sigma}{\sigma\tau}+\frac{n-1}{\tau}+o(1).$
Hereby we consider $\tau\geq 1,\sigma\geq 1$ absolute so that the expression is minimized upon all pairs $\geq 1$. The right hand side clearly decays in $\sigma$, hence we can put $\sigma=1$ and get
$\kappa(n,\xi,X)\leq\frac{1}{\tau}((\tau-1)w_{k}(\xi)+n-1)+o(1)=w_{k}(\xi)+\frac{n-1-w_{k}(\xi)}{\tau}+o(1).$
If $n-1\geq w_{k}(\xi)$ then the bound decreases in $\tau$, hence we may put $\tau=1$ for all large $X$ to get the bound $n-1$. If otherwise $n-1\leq w_{k}(\xi)$ then we have the bound $w_{k}(\xi)$. Since $X$ was arbitrary, the stronger claim (14) follows as $X\to\infty$ and (13) is implied by (7).
6. Proof of Theorem 2.2
6.1. Another lemma
We first verify
Lemma 6.1.
Assume for given $n,\xi$ and for all large $X$, the polynomial $P_{X}$ is separable. Then (11) holds, in fact $\overline{\kappa}_{n}(\xi)\leq n-1$. More generally for every $X$ with this property
$\kappa(n,\xi,X)\leq n-1.$
Proof.
We may apply Lemma 4.2 to $P_{X}$. Proceeding as in the proof of Theorem 3.1 (or Theorem 3.6) with $P=Q$ and thus $\sigma=\tau$, the polynomial $P_{X}$ of degree $m\leq n$ has a root $\alpha$ with
$|\xi-\alpha|\ll_{n}H(\alpha)^{-1}X^{-\frac{\sigma w-m+1}{\sigma}}.$
In other words
$w^{\ast}(n,\xi,X)\geq\frac{\sigma w-m+1}{\sigma}\geq w-\frac{n-1}{\sigma}\geq w-n+1.$
The latter claim is now immediate, and the first follows by (7) if the separability hypothesis holds for all large $X$. ∎
We remark that thanks to this lemma we can improve the bound of Corollary 3.2 for very small $n$.
Corollary 6.2.
We have
$0=\dim_{H}(\Omega_{2})\leq\dim_{H}(\Gamma_{2})\leq\frac{2}{3}$
and
$0=\dim_{H}(\Omega_{3})\leq\dim_{H}(\Gamma_{3})\leq\frac{1}{2}.$
From the argument of Corollary 3.2 we would get the bounds $1$ and $2/3$ respectively.
Proof.
All but the upper bounds for $\dim_{H}(\Gamma_{2}),\dim_{H}(\Gamma_{3})$ are clear by Theorem 2.1. If for all large $X$ the best approximation $P_{X}$ is separable, then Lemma 6.1 implies $\overline{\kappa}_{n}(\xi)\leq n-1$, so we can assume the opposite. But for $n\leq 3$, this means that $P_{X}$ decomposes into linear factors for certain arbitrarily large $X$. This implies $w_{1}(\xi)\geq n$ by Wirsing’s Lemma 4.1, and the estimates follow from (15) applied to $m=1$. ∎
6.2. Proof of Theorem 2.2
If $\widehat{w}_{n}(\xi)=n$ then by (6) trivially
(35) $\widehat{w}_{n}^{\ast}(\xi)\geq 1=n-n+1=\widehat{w}_{n}(\xi)-n+1.$
Now assume $\widehat{w}_{n}(\xi)>2n-3$. By Theorem 2.1 we may assume $n\geq 3$ (in fact $n\geq 4)$. Assume $P_{X}$ is not separable for some large $X$. Then it contains a square of a polynomial as a factor. But by $n\geq 3$ this means that any irreducible factor has degree at most $\max\{\lfloor n/2\rfloor,n-2\}=n-2$. Thus Wirsing’s Lemma 4.1 and (4) imply
$w_{n-2}(\xi)\geq\widehat{w}_{n}(\xi)\geq n.$
However by Theorem 4.5 this implies
$\widehat{w}_{n}(\xi)=\min\{\widehat{w}_{n}(\xi),w_{n-2}(\xi)\}\leq n+(n-2)-1=2n-3.$
Thus assuming the reverse inequality $\widehat{w}_{n}(\xi)>2n-3$ all best approximation polynomials of large norm are separable. Thus the claim follows from Lemma 6.1.
Now assume $\widehat{w}_{n}^{\ast}(\xi)\geq n-2$. Then the claim is obviously true if $\widehat{w}_{n}(\xi)\leq(n-2)+(n-1)=2n-3$. If this is false then the implication follows from (ii).
Assume (iv) holds. If $\lambda_{n}(\xi)\leq 1/(n-2)$ then by the main result of [15]
$\widehat{w}_{n}^{\ast}(\xi)\geq\frac{1}{\lambda_{n}(\xi)}\geq n-2$
which is hypothesis (iii). If $\lambda_{n}(\xi)>1$, then it follows from  Theorem 4.6 $\widehat{w}_{n}(\xi)=n$, so we reduced it to (i). This also contains the implication from (v).
Finally for the conclusion from (vi), combining (3.3) and (7) yields
$\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq\overline{\kappa}_{n}(\xi)\leq w_{n}(\xi)-\frac{w_{n}(\xi)}{w_{n}(\xi)-n+1}.$
A short calculation shows that the right hand side is at most $n-1$ if $w_{n}(\xi)$ is bounded as in (vi).
7. Proof of Theorem 2.1
7.1. Proof for $n\leq 3$
Let $n\leq 3$. Then if $P_{X}$ is separable, it decomposes into linear factors. If this happens for certain arbitrarily large $X$, by Wirsing’s Lemma 4.1 and (4) this implies that
$w_{1}(\xi)\geq\widehat{w}_{n}(\xi)\geq n.$
We conclude via Theorem 2.2, (v). If otherwise there is no such sequence of $X$ that tends to infinity, then (11) holds as well by Lemma 6.1.
7.2. Towards the proof for $n\in\{4,5\}$. A preparatory lemma
The cases $n=4,n=5$ are considerably more complicated. We prove a general auxiliary fact for each $n$.
Let us first fix some notation. Let $X>1$ be any large number and $P=P_{X}$ the according best approximation polynomial. Write
$P=\prod_{i=1}^{\ell}Q_{i}^{\alpha_{i}},\qquad\alpha_{i}\in\mathbb{N}$
for its factorization into irreducible polynomials. Let further
$d_{i}=\deg Q_{i}\geq 1.$
Note that
(36) $\sum_{i=1}^{\ell}\alpha_{i}d_{i}\leq n.$
Let $H_{i}:=H(Q_{i})$ and define $\gamma_{i}>0$ by
$|Q_{i}(\xi)|=H_{i}^{-\gamma_{i}}.$
Define $\beta_{i}>0$, $1\leq i\leq\ell$, by
$H_{i}=X^{\beta_{i}}.$
Gelfond’s Lemma 4.3 implies $\prod X^{\alpha_{i}\beta_{i}}=\prod H_{i}^{\alpha_{i}}\ll_{n}H_{P}\leq X$ and thus
(37) $\sum_{i=1}^{\ell}\alpha_{i}\beta_{i}\leq 1+o(1).$
Now since
$|P(\xi)|=\prod_{i=1}^{\ell}|Q_{i}(\xi)|^{\alpha_{i}}=\prod_{i=1}^{\ell}H_{i}^{-\alpha_{i}\gamma_{i}}=X^{-\sum\alpha_{i}\beta_{i}\gamma_{i}}$
we get
(38) $w(n,\xi,X)\leq\sum_{i=1}^{\ell}\alpha_{i}\beta_{i}\gamma_{i}+o(1).$
On the other hand, considering the separable polynomial
$Q:=\prod_{i=1}^{\ell}Q_{i}$
which satisfies
$\deg Q=\sum_{i=1}^{\ell}d_{i},\qquad|Q(\xi)|=\prod_{i=1}^{\ell}|Q_{i}(\xi)|=\prod_{i=1}^{\ell}H_{i}^{-\gamma_{i}},$
we get with Lemma 4.2 and using $H_{Q}\asymp_{n}\prod_{i=1}^{\ell}H_{i}$ by Gelfond’s Lemma 4.3 that it has a root $\alpha$ with
$|\alpha-\xi|\ll_{n}H(\alpha)^{-1}\cdot\prod_{i=1}^{\ell}H_{i}^{-\gamma_{i}+d_{i}-1}=H(\alpha)^{-1}\cdot X^{-\sum_{i=1}^{\ell}\beta_{i}(\gamma_{i}-d_{i}+1)}$
thus
(39) $w^{\ast}(n,\xi,X)\geq\sum_{i=1}^{\ell}\beta_{i}\cdot(\gamma_{i}-d_{i}+1).$
Comparing (38), (39) we get
$\kappa(n,\xi,X)\leq\sum_{i=1}^{\ell}\alpha_{i}\beta_{i}\gamma_{i}-\sum_{i=1}^{\ell}\beta_{i}\cdot(\gamma_{i}-d_{i}+1),$
We have thus proved the following lemma.
Lemma 7.1.
For any large $X$ with induced parameters $\ell,d_{i},\alpha_{i},\beta_{i},\gamma_{i}$ we have
$\kappa(n,\xi,X)\leq\sum_{i=1}^{\ell}(\alpha_{i}-1)\beta_{i}\gamma_{i}+\beta_{i}(d_{i}-1)$
Thus if we have
(40) $\sum_{i=1}^{\ell}(\alpha_{i}-1)\beta_{i}\gamma_{i}+\beta_{i}(d_{i}-1)\leq-1+\sum_{i=1}^{\ell}\alpha_{i}d_{i}\leq n-1$
then
(41) $\kappa(n,\xi,X)\leq n-1.$
Remark 2.
A similar argument can be applied to $Q:=\prod_{j\in J}Q_{j}$ for any subset $J\subseteq\{1,2,\ldots,\ell\}$. However it turns out one may restrict to the full set $J$.
We prove the cases $n=4,n=5$ with Lemma 7.1 and Theorems 4.5, 4.6. By Lemma 6.1 we only have to consider parameters $X$ for which $P_{X}$ is not separable. Moreover, we can assume that for all large $X$ it does not compose into linear factors, as otherwise by Wirsing’s Lemma 4.1 we have $w_{1}(\xi)\geq\widehat{w}_{n}(\xi)\geq n$ and the claim follows from Theorem 2.2, (v).
We may assume $P_{X}$ has exact degree $n$. Otherwise we could multiply it with the according power of the variable $T$ to get $\tilde{P}(T)=T^{t}P_{X}$ for $t=n-\ \deg P_{X}$, which keeps the height unaffected $H_{\tilde{P}}=H_{P}$ and decreases its evaluation at $\xi$. i.e. $|\tilde{P}(\xi)|<|P(\xi)|$ (as we may assume $\xi\in(0,1)$), contradicting the definition of $P_{X}$.
We will use the notation of the present § 7.2 in the sequel.
7.3. Proof for $n=4$
First let $n=4$. In view of the above restrictions, for $X$ any large parameter, there are two remaining cases.
Case A: We have
$\ell=1,\qquad P_{X}=Q_{1}^{2}Q_{2},\qquad d_{1}=1,\;d_{2}=2,$
that is $Q_{1}$ is linear and $Q_{2}$ irreducible quadratic. Then (37) gives
(42) $2\beta_{1}+\beta_{2}\leq 1,$
in particular
$\beta_{1}\leq 1/2.$
Then the criterion (40) from Lemma 7.1 yields the sufficient condition
$\beta_{1}\gamma_{1}+\beta_{2}\leq n-1=3.$
Using (42) this can be relaxed to $(\gamma_{1}-2)\beta_{1}+1\leq 3$ or equivalently
$\gamma_{1}\leq 2+\frac{2}{\beta_{1}}.$
In particular $\gamma_{1}\leq 6$ suffices to have (41) for the corresponding $X$ of Case A. Note that by definition of $\widehat{w}_{n}(\xi)$ this further implies
(43) $\widehat{w}_{n}(\xi)-w^{\ast}(n,\xi,X)\leq n-1+o(1),\qquad X\to\infty,$
for these $X$.
However, if $\gamma_{1}>6$ for arbitrarily large $X$ inducing Case A then $w_{1}(\xi)>6>4=n$, implying (11) by Theorem 2.2, (v). In particular we again have (43). So (43) holds for all $X$ inducing Case A.
Case B: We have
$P_{X}=Q_{1}^{2},\qquad d_{1}=2.$
Then $Q_{1}$ is quadratic irreducible and
$\ell=1,\qquad\alpha_{1}=\alpha=2,\qquad 0<\beta_{1}=\beta\leq\frac{1}{2},\qquad d_{1}=d=2,\qquad\gamma_{1}=\gamma.$
Now the sufficient criterion (40) of Lemma 7.1 becomes
$\beta(\gamma+1)\leq 3$
or
(44) $\gamma\leq\frac{3}{\beta}-1.$
In particular
$\gamma\leq 5$
suffices to prove
$\kappa(4,\xi,X)=w(4,\xi,X)-w^{\ast}(4,\xi,X)\leq 3=n-1$
for $X$ in question. Assume otherwise
$\gamma>5=n+d-1.$
If this happens for arbitrarily large $X$, then by Theorem 4.5 we have
$\widehat{w}_{n}(\xi)=\min\{\widehat{w}_{n}(\xi),w_{d}(\xi)\}\leq n+d-1=n+1.$
So it suffices to show
(45) $w^{\ast}(n,\xi,X)\geq 2-o(1),\qquad X\to\infty,$
for $X$ inducing Case B, to deduce that for such $X$ again (43) holds. Assume this is true. As we have observed (43) holds for any large $X$ inducing Case A as well, we can pass to lower limit to indeed deduce $\widehat{w}_{n}(\xi)-\widehat{w}_{n}^{\ast}(\xi)\leq n-1$, i.e. (11).
However, for $X$ inducing Case B, by Feldman’s Lemma 4.2 the quadratic $Q_{1}=Q$ has a root $\alpha$ with
$|\alpha-\xi|\ll H_{Q}^{-\gamma}=H(\alpha)^{-1}H_{Q}^{-(\gamma-1)}=H(\alpha)^{-1}X^{-\beta(\gamma-1)}.$
Hence if
$\beta(\gamma-1)\geq 2$
or equivalently
(46) $\gamma\geq 1+\frac{2}{\beta}$
we have (45) and are done. So combining the criteria (44), (46), the only problematic case is
$\frac{3}{\beta}-1<\gamma<1+\frac{2}{\beta}.$
But this implies $\beta>1/2$, contradiction.
Remark 3.
We stress again that Theorem 4.5 and Theorem 4.6 were vital ingredients of the proof of both Cases A,B. Indeed, the stronger claim $\kappa(4,\xi,X)\leq n-1$ is false in either case in general, by a generalization of Theorem 3.3. The same will be true likewise for $n=5$ below.
7.4. Proof for $n=5$
We now prove the claim for $n=5$. Here by the initial argument as for $n=4$ we have now four remaining cases:
Case A:
$P_{X}=Q_{1}^{2}Q_{2},\qquad d_{1}=1,\quad d_{2}=3.$
Then
$2\beta_{1}+\beta_{2}\leq 1$
in particular
$\beta_{1}\leq\frac{1}{2}$
and (40) becomes
$\beta_{1}\gamma_{1}+2\beta_{2}\leq n-1=4$
so sufficient condition is
$\gamma_{1}\leq\frac{2}{\beta_{1}}+4.$
Hence
$\gamma_{1}\leq 6$
suffices. But if for arbitrarily large such $X$
$\gamma_{1}>6>5=n$
then $w_{1}(\xi)\geq n$ and we conclude (43) for $X$ in question by the same arguments as for $n=4$.
Case B:
$P_{X}=Q_{1}^{3}Q_{2},\qquad d_{1}=1,\;d_{2}=2$
Then
$3\beta_{1}+\beta_{2}\leq 1$
thus
$\beta_{1}\leq\frac{1}{3}$
and (40) leads to
$2\beta_{1}\gamma_{1}+2\beta_{1}\leq 4=n-1,$
or
$\beta_{1}(\gamma_{1}+1)\leq 2,\qquad\gamma_{1}\leq\frac{2}{\beta_{1}}-1,$
so
$\gamma_{1}\leq 5$
suffices. But if for arbitrarily large $X$ we have
$\gamma_{1}>5=n$
then again $w_{1}(\xi)\geq n$ and we again conclude that (43) holds.
Case C:
$P_{X}=Q_{1}^{2}Q_{2}Q_{3},\qquad d_{1}=d_{2}=1,\;d_{3}=2.$
Then
$2\beta_{1}+\beta_{2}+\beta_{3}\leq 1$
in particular
$\beta_{1}\leq\frac{1}{2}$
and (40) yields the criterion
$\beta_{1}\gamma_{1}+\beta_{3}\leq 4=n-1,$
hence as $\beta_{2}\geq 0$ we get
$\beta_{1}(\gamma_{1}-2)+1\leq 4,\qquad\gamma_{1}\leq\frac{3}{\beta_{1}}+2$
as sufficient criterion. Now if conversely for arbitrarily large $X$ we have
$\gamma_{1}>\frac{3}{\beta_{1}}+2\geq 8,$
then again $w_{1}(\xi)\geq 8>n$ and we conclude again that (43) holds.
Case D:
$P_{X}=Q_{1}^{2}Q_{2},\qquad d_{1}=2,\;d_{2}=1.$
Then again
$2\beta_{1}+\beta_{2}\leq 1$
in particular
$\beta_{1}\leq\frac{1}{2}$
and (40) becomes
$\beta_{1}\gamma_{1}+\beta_{1}=\beta_{1}(\gamma_{1}+1)\leq n-1=4$
so sufficient is
$\gamma_{1}\leq\frac{4}{\beta_{1}}-1.$
Hence
$\gamma_{1}\leq 7$
suffices to prove
$w(5,\xi,X)-w^{\ast}(5,\xi,X)\leq n-1$
for all such $X$. But if for arbitrarily large $X$
$\gamma_{1}>7>6=n+d_{1}-1$
holds, then by Theorem 4.5 we have
$\widehat{w}_{n}(\xi)=\min\{\widehat{w}_{n}(\xi),w_{d_{1}}(\xi)\}\leq n+d_{1}-1=n+1.$
So we only have to prove that
$w^{\ast}(5,\xi,X)\geq 2-o(1),\qquad X\to\infty,$
for $X$ inducing Case D, to conclude first (43) and finally (11) by the same arguments as for $n=4$. By Feldman’s Lemma 4.2, for this again with the same argument as in Case B of $n=4$ it suffices to have
$\gamma_{1}\geq 1+\frac{2}{\beta_{1}}.$
So the only bad case is
$\frac{4}{\beta_{1}}-1<\gamma_{1}<1+\frac{2}{\beta_{1}}.$
But this implies $\beta_{1}>1$, contradiction.
8. Proof of Theorem 3.3
We only need to prove (16), then (17) is just the case $k=n$ and (18) follows in turn for $w_{1}(\xi)=\infty$. Let $Q(T)=aT-b$ with $a>0$ be very small at $\xi$, say
$|Q(\xi)|=H_{Q}^{-\lambda}$
for some $\lambda>n+k-1$. Clearly we can assume $|b|\asymp_{\xi}a$ so this implies
(47) $|\xi-b/a|=\frac{|Q(\xi)|}{a}\asymp_{\xi}H_{Q}^{-\lambda-1}.$
Consider the parameter
$X=H_{Q^{k}}\asymp_{n}H_{Q}^{k},$
where we used Gelfond’s Lemma 4.3. Consider $R=Q^{k}$, which has degree $k\leq n$ and $H_{R}=X$. Moreover
$|R(\xi)|=H_{Q}^{-k\lambda}\ll_{n}X^{-\lambda}.$
Hence
$w(n,\xi,X)\geq\lambda-o(1),$
as $H_{Q}\to\infty$.
On the other hand, any other algebraic number $\eta$ of degree at most $n$ and height at most $X$ has distance from $b/a$ at least
$|\eta-b/a|\gg_{n}X^{-1}H_{Q}^{-n}\gg_{n}X^{-1-n/k}$
by Liouville inequality Theorem 4.4. So by $\lambda>n+k-1$, this implies by (47) and triangle inequality for large enough $X$
$|\xi-\eta|\geq|b/a-\eta|-|\xi-b/a|\geq\frac{1}{2}|b/a-\eta|\gg_{n}X^{-1-n/k}.$
So the smallest value comes from $b/a$ the root of $Q$, for which we use (47) and estimating its right hand side as $\gg H(b/a)^{-1}X^{-\lambda/k-o(1)}$ we get
$w^{\ast}(n,\xi,X)\leq\frac{\lambda}{k}+o(1).$
Combining the two bounds we get
$\kappa(n,\xi,X)\geq\lambda-\frac{\lambda}{k}-o(1)=\left(1-\frac{1}{k}\right)\lambda-o(1).$
As $\lambda$ can be chosen arbitrarily close to $w_{1}(\xi)$ and by assumption $w_{1}(\xi)>n+k-1$, we can find infinitely many such $Q$ and thus arbitrarily large $X$, and the claim follows.
9. Proof of Theorem 3.5
We prove an auxiliary lemma for approximation to a single number using Minkowski’s Second Convex Body Theorem.
Lemma 9.1.
Let $P(T)=cT+d$ be a linear integer polynomial best approximation for $n=1$ (essentially this means $-d/c$ is a convergent to $\xi$). Let $\lambda\geq 1$ be determined by
$|P(\xi)|=H_{P}^{-\lambda}.$
Let $R(T)=aT+b$ be the linear integer polynomial that minimizes $|Q(\xi)|$ among all linear integer polynomials $Q(T)$ of height $H_{Q}\leq C\cdot H_{P}^{\lambda}$ for fixed $C\in(0,\frac{1}{2})$ that are not a scalar multiple of $P$. Then
$|R(\xi)|\asymp H_{P}^{-1}.$
The implied constants are absolute.
Proof.
By well known estimates for continued fractions, for
$|R(\xi)|\ll H_{P}^{-1}$
it suffices to take $R(T)=Q(T)=aT+b$ the best approximation polynomial for the parameter $Y=CH_{P}$. This is clearly not a multiple of $P$ as $C<1$ and $P$ has a primitive coefficient vector since it is a best approximation polynomial. Moreover as $\lambda\geq 1$ clearly $H_{Q}\leq C\cdot H_{P}\leq C\cdot H_{P}^{\lambda}$, and by Dirichlet’s Theorem and definition of $R$ indeed
$H_{R}\leq C\cdot H_{P},\qquad|R(\xi)|\leq|Q(\xi)|\leq(C\cdot H_{P})^{-1}\ll H_{P}^{-1}.$
(Usually $R$ corresponds to the convergent $-b/a$ of $\xi$ preceding $-d/c$ in the continued fraction algorithm).
For the reverse inequality, assume
$H_{R}\leq C\cdot H_{P},\qquad|R(\xi)|\leq D\cdot H_{P}^{-1}$
for some $D\in(0,\frac{1}{2})$. We use Minkowski’s Second Convex Body Theorem. Take the parameter
$Z=H_{P}^{(\lambda+1)/2}$
and consider the successive minima for the convex body $K_{Z}$ of points $(a_{0},a_{1})\in\mathbb{R}^{2}$ satisfying
$-Z\leq|a_{1}|\leq Z,\qquad-Z^{-1}\leq|a_{0}+a_{1}\xi|\leq Z^{-1}$
of area $4$ and the lattice $\mathbb{Z}^{2}$. Denote them by $\lambda_{1}(K_{Z},\mathbb{Z}^{2})$ and $\lambda_{2}(K_{Z},\mathbb{Z}^{2})$. Since $P,R$ are linearly independent they induce at most the first and second successive minimum respectively. It can be checked that $P$ induces
$\lambda_{1}(K_{Z},\mathbb{Z}^{2})\leq Z^{-(\lambda-1)/(\lambda+1)}.$
Indeed
$Z^{-1-(\lambda-1)/(\lambda+1)}=H_{P}^{-\lambda}=|P(\xi)|$
and
$Z^{1-(w-1)/(w+1)}=H_{P}.$
This shows that the integer point corresponding to $P$ lies in $Z^{-(w-1)/(w+1)}K_{Z}$. Similarly $R$ induces
$\lambda_{2}(K_{Z},\mathbb{Z}^{2})\leq E\cdot Z^{(\lambda-1)/(\lambda+1)}$
for $E:=\max\{C,D\}\in(0,\frac{1}{2})$. Indeed
$E\cdot Z^{-1+(\lambda-1)/(\lambda+1)}\geq D\cdot Z^{-1+(\lambda-1)/(\lambda+1)}=H_{P}^{-1}\geq|R(\xi)|$
and
$E\cdot Z^{1-(\lambda-1)/(\lambda+1)}\geq C\cdot Z^{1-(\lambda-1)/(\lambda+1)}=C\cdot H_{P}^{\lambda}\geq H_{R},$
which again shows $R$ lies in $E\cdot Z^{(\lambda-1)/(\lambda+1)}K_{Z}$. Hence the product satisfies
$\lambda_{1}(K_{Z},\mathbb{Z}^{2})\lambda_{2}(K_{Z},\mathbb{Z}^{2})\leq E.$
However, since $E<1/2$ this contradicts Minkowski’s Second Convex Body Theorem stating the product is bounded from below by $\lambda_{1}(K_{Z},\mathbb{Z}^{2})\lambda_{2}(K_{Z},\mathbb{Z}^{2})\geq 1/2$. ∎
Now we prove the theorem. Let $P,R$ be as in Lemma 9.1 and assume $\lambda>3$. By Lemma 9.1 they satisfy
$|P(\xi)|=H_{P}^{-\lambda},\quad|R(\xi)|\asymp H_{P}^{-1}.$
Hence the polynomials
$V_{1}:=P^{2},\qquad V_{2}:=PR$
are quadratic and satisfy
(48) $\max_{i=1,2}H_{V_{i}}\ll H_{P}^{2},\qquad|V_{1}(\xi)|=H_{P}^{-2\lambda},\quad|V_{2}(\xi)|\asymp H_{P}^{-\lambda-1}.$
Take the parameter
$X=H_{P}^{\lambda-1-\varepsilon}$
for small $\varepsilon\in(0,(\lambda-3)/2)$ and consider all linear or quadratic integer polynomials $V_{3}$ of height at most $H_{V_{3}}\leq X$. Clearly any polynomial in $\rm{span}\{V_{1}\}$ is not separable, hence for our exponent we can restrict to the complementary set of polynomials. We show that any such $V_{3}\notin\rm{span}\{V_{1}\}$ with $H_{V_{3}}\leq X$ satisfies
(49) $|V_{3}(\xi)|\gg H_{P}^{-\lambda-1}.$
If this is true then
$w_{sep}(2,\xi,X)\leq\frac{\lambda+1+o(1)}{\lambda-1-\varepsilon}=1+\frac{2+\varepsilon+o(1)}{\lambda-1-\varepsilon}.$
The claim (22) of the theorem follows with $H_{P}\to\infty$ as we may take $\varepsilon$ arbitrarily small and $\lambda$ arbitrarily close to $w_{1}(\xi)$. The latter specialization for Liouville numbers follows as then the upper bound becomes $1$, while by considering linear polynomials only we get the reverse inequality, to conclude
$1\geq\widehat{w}_{2,sep}(\xi)\geq\widehat{w}_{2,irr}(\xi)\geq\widehat{w}_{1,irr}(\xi)=\widehat{w}_{1}(\xi)\geq 1$
So the inequalities must be equalities. To prove (49), we distinguish two cases.
Case 1: $V_{3}\in\rm{span}\{V_{1},V_{2}\}\setminus\rm{span}\{V_{1}\}$. This means $V_{3}=PS$ with linear $S\notin\rm{span}\{P\}$. Moreover by Gelfond’s Lemma 4.3 we have $H_{S}\ll H_{V_{3}}/H_{P}\leq X/H_{P}=H_{P}^{w-2-\varepsilon}$, hence we may assume $H_{S}<C\cdot H_{P}^{\lambda}$ if we assume $H_{P}$ is large enough. By Lemma 9.1 we conclude
$|S(\xi)|\gg H_{P}^{-1}$
and thus
$|V_{3}(\xi)|=|P(\xi)|\cdot|S(\xi)|\gg H_{P}^{-\lambda-1},$
so (49) holds.
Case 2: $V_{3}\notin\rm{span}\{V_{1},V_{2}\}$. So assume contrary to (49) that some linear or quadratic integer polynomial $V_{3}$ (in fact it is not hard to show that $V_{3}$ must be irreducible quadratic) outside the span of $\{V_{1},V_{2}\}$ satisfies the estimates
(50) $H_{V_{3}}\leq H_{P}^{\lambda-1-\varepsilon},\qquad|V_{3}(\xi)|\leq H_{P}^{-\lambda-1}.$
We proceed similarly as in the proof of Lemma 9.1. Take the parameter
$Q=H_{P}^{\frac{4}{3}\lambda}$
and define the convex body $\mathcal{K}_{Q}$ of $(a_{0},a_{1},a_{2})\in\mathbb{R}^{3}$ satisfying
$\max_{i=1,2}|a_{i}|\leq Q^{1/2},\qquad|a_{0}+\xi a_{1}+\xi^{2}a_{2}|\leq Q^{-1}.$
Note $\mathcal{K}_{Q}$ has constant volume $8$ independent of $Q$. Then as $V_{1},V_{2},V_{3}$ are linearly independent, they induce at most the first, second and third successive minimum with respect to $\mathcal{K}_{Q}$ and the lattice $\Lambda=\mathbb{Z}^{3}$. Denote them by $\lambda_{1}(\mathcal{K}_{Q},\mathbb{Z}^{3}),\lambda_{2}(\mathcal{K}_{Q},\mathbb{Z}^{3}),\lambda_{3}(\mathcal{K}_{Q},\mathbb{Z}^{3})$. However, it can be checked via (48) that $V_{1}$ induces
$\lambda_{1}(\mathcal{K}_{Q},\mathbb{Z}^{3})\leq Q^{\chi+o(1)},\qquad\chi:=-\frac{\lambda-3}{2\lambda}<0$
since
$Q^{1/2+\chi}=Q^{3/(2\lambda)}=H_{P}^{2}\asymp H_{V_{1}},\qquad Q^{-1+\chi}=Q^{-3/2+3/(2\lambda)}>Q^{-3/2}=H_{P}^{-2\lambda}=|V_{1}(\xi)|,$
so indeed the integer point corresponding to $V_{1}$ lies in $Q^{\chi+o(1)}\cdot\mathcal{K}_{Q}$. Similarly via (48) and (50) the polynomials $V_{2},V_{3}$ induce
$\lambda_{3}(\mathcal{K}_{Q},\mathbb{Z}^{3})\leq\lambda_{2}(\mathcal{K}_{Q},\mathbb{Z}^{3})\ll Q^{\eta+o(1)},\qquad\eta:=\frac{1}{2}\cdot\frac{\lambda-3-2\varepsilon}{2\lambda-\varepsilon}>0.$
Indeed by $\varepsilon>0$ one can verify
$Q^{1/2+\eta}\geq Q^{(\lambda-1)\frac{3}{4\lambda}}=H_{P}^{\lambda-1}\gg\max\{H_{V_{2}},H_{V_{3}}\}$
and again by $\varepsilon>0$
$Q^{-1+\eta}>Q^{-(\lambda+1)\frac{3}{4\lambda}}=H_{P}^{-\lambda-1}\geq\max\{|V_{2}(\xi)|,|V_{3}(\xi)|\},$
showing together that the points corresponding to $V_{i},i=2,3$, lie in $Q^{\eta+o(1)}\cdot\mathcal{K}_{Q}$. Since $\varepsilon>0$ we have $\chi+2\eta<0$, so the product of the successive minima
$\lambda_{1}(\mathcal{K}_{Q},\mathbb{Z}^{3})\lambda_{2}(\mathcal{K}_{Q},\mathbb{Z}^{3})\lambda_{3}(\mathcal{K}_{Q},\mathbb{Z}^{3})\ll Q^{\chi+2\eta+o(1)}$
tends to 00 as $H_{P}\to\infty$. This contradicts Minkowski’s Second Convex Body Theorem stating it is bounded from below by an absolute positive constant.
10. Proof of Theorem 3.6
For (25), note that the weaker estimate
$\underline{\kappa}_{2}(\xi)\leq w_{2}(\xi)-w_{2}^{\ast}(\xi)\leq w_{2}(\xi)-\widehat{w}_{2}(\xi)^{2}+\widehat{w}_{2}(\xi)\leq w_{2}(\xi)-2$
follows directly from combining [14, Theorem 2] (see also [9]) and (8). For the stronger claim, it follows more precisely from the proof of [14] that there are infinitely many best approximations satisfying
$|P^{\prime}(\xi)|\geq|P(\xi)|\cdot H_{P}^{\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi)+1}$
which implies $P$ has a root $\alpha$ with
$|\xi-\alpha|\ll H(\alpha)^{-1}\cdot H_{P}^{-(\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi))+o(1)}.$
Moreover, a result of Jarník [11, 12] implies that $P$ remains a best approximations up to a parameter of size $X\geq H_{P}^{\sigma-o(1)}$ for $\sigma=\widehat{w}_{2}(\xi)-1$. Indeed this happens whenever $P$ is linearly independent with its preceding and successive minimal polynomial, which occurs for our $P$ above. Hence
$|\xi-\alpha|\ll H(\alpha)^{-1}H_{P}^{-(\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi))+o(1)}=H(\alpha)^{-1}X^{-\frac{\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi)}{\sigma}+o(1)}.$
In other words
$w^{\ast}(2,\xi,X)\geq\frac{\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi)}{\sigma}-o(1),$
hence since clearly $w(2,\xi,X)\leq w_{2}(\xi)/\sigma+o(1)$ we get
$\kappa(2,\xi,X)\leq w(2,\xi,X)-\frac{\widehat{w}_{2}(\xi)^{2}-\widehat{w}_{2}(\xi)}{\sigma}+o(1)\leq\frac{w_{2}(\xi)}{\widehat{w}_{2}(\xi)-1}-\widehat{w}_{2}(\xi)+o(1).$
The claim follows as there are arbitrarily large such $X$.
For (26), we again use $P,X$ as above and for simplicity write $w=w(2,\xi,X)$. By Lemma 4.2 below, $P$ has a root $\alpha$ satisfying
$|\xi-\alpha|\ll|P(\xi)|=X^{-w}=H(\alpha)^{-1}X^{-w}H_{P}=H(\alpha)^{-1}X^{-\frac{\sigma w-1}{\sigma}}.$
In other words
$w^{\ast}(2,\xi,X)\geq w-\frac{1}{\sigma}-o(1).$
Rearranging, for such $X$ we have
$\kappa(2,\xi,X)=w(2,\xi,X)-w^{\ast}(2,\xi,X)\leq\frac{1}{\sigma}+o(1)\leq\frac{1}{\widehat{w}_{2}(\xi)-1}+o(1),$
so the claim follows.
The author thanks Yann Bugeaud for helpful remarks, especially for bringing to my attention a question that led to Theorem 3.5.